# Dream, Lift, Animate: From Single Images to Animatable Gaussian Avatars

### v1

@@ Abstract
We introduce Dream, Lift, Animate (DLA), a novel framework that reconstructs animatable 3D human avatars from a single image. This is achieved by leveraging multi-view generation, 3D Gaussian lifting, and pose-aware UV-space mapping of 3D Gaussians. Given an image, we first dream plausible multi-views using a video diffusion model, capturing rich geometric and appearance details. These views are then lifted into unstructured 3D Gaussians. To enable animation, we propose a transformer-based encoder that models global spatial relationships and projects these Gaussians into a structured latent representation aligned with the UV space of a parametric body model. This latent code is decoded into UV-space Gaussians that can be animated via body-driven deformation and rendered conditioned on pose and viewpoint. By anchoring Gaussians to the UV manifold, our method ensures consistency during animation while preserving fine visual details. DLA enables real-time rendering and intuitive editing without requiring post-processing. Our method outperforms state-of-the-art approaches on ActorsHQ and 4D-Dress datasets in both perceptual quality and photometric accuracy. By combining the generative strengths of video diffusion models with a pose-aware UV-space Gaussian mapping, DLA bridges the gap between unstructured 3D representations and high-fidelity, animation-ready avatars.
https://research.nvidia.com/labs/dair/dream-lift-animate
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Figure 1: We propose Dream, Lift, Animate, a novel framework to reconstruct high-fidelity, animatable 3D human avatars from a single image by generating multi-view images, lifting them to 3D Gaussians, and mapping them to a pose-aware UV space (Fig. 2). Our approach enables realistic animation and outperforms prior methods in visual quality (Fig. 6 and Tbl. 1).
@@ 1 Introduction
Figure 2: Overview of the proposed Dream, Lift, Animate (DLA) framework for reconstructing animatable 3D human avatars from a single image. In the Dream stage, we synthesize novel views from the input using a diffusion-based generator. In the Lift stage (Fig. 3), we project the multi-view images into a set of unstructured 3D Gaussians in the pose space using a learned Gaussian reconstruction model $\mathcal{G}$. Subsequently, we learn a transformer encoder $\mathcal{F}$ to map 3D Gaussians to a structured latent code $\mathbf{Z}$ in the UV space of a parametric body model. In the Animate stage (Fig. 4), we decode the avatar code into a pose- and view-aware Gaussian parameter map $\mathbf{F}$. This structured representation enables realistic animation and rendering via deformation with a body model.
@@
Existing methods typically fall short in simultaneously addressing these challenges. Recent approaches ho2024sith ; xue2024human3diffusion ; he2024magicman leverage the powerful generative capabilities of multi-view diffusion models Rombach2022StableDiffusion to hallucinate missing appearance information; however, these models inherently produce inconsistencies across generated views, leading to avatars that appear overly smoothed, lack critical details, or introduce visually jarring artifacts when animated. Conversely, methods relying on template-based rigging approaches, such as automatically transferring skinning weights from canonical human body models (e.g., SMPL loper2015smpl ), typically require careful fitting procedures zhang2023sifu ; huang2022elicit ; huang2024tech ; svitov2023dinar . While powerful, these methods can still encounter challenges when handling non-standard poses, complex clothing, and significant self-occlusions frequently encountered in unconstrained, real-world scenarios, thus affecting their robustness and broader applicability.
@@
To overcome these limitations, we propose Dream, Lift, Animate (DLA), a novel framework designed specifically for reconstructing high-quality, animatable 3D human avatars from a single image. Our method addresses the core reconstruction problem by decomposing it into three complementary stages, each carefully designed to handle limitations inherent in the preceding steps. Specifically, we first leverage a pretrained video diffusion model wang2024unianimate to dream plausible multi-view observations of the subject, effectively hallucinating realistic geometric and appearance details even in regions unseen from the input viewpoint. Despite the significant progress made by recent diffusion-based approaches, their inherent view-to-view inconsistencies prevent direct avatar reconstruction xue2024human3diffusion . To resolve this, we next lift these generated views into an intermediate, unstructured 3D representation based on 3D Gaussian primitives kerbl20233d in the input pose space, effectively aggregating appearance and geometry cues from the multi-view observations. Subsequently, to enable animation, we map these unstructured Gaussians into a structured UV-space representation. For this, we use a transformer-based encoder zhang3dshape2vecset that models global spatial relationships among the unstructured 3D Gaussians and projects them into a structured latent avatar representation, which aligns with the UV space of the SMPL-X body model pavlakos2019smplx . This learned mapping also allows our model to reconcile inconsistencies introduced during the initial multi-view generation seamlessly. Finally, in the animate stage, we present a pose- and view-aware Gaussian parameter decoder that converts this latent avatar representation into a UV-space map of Gaussian parameters, enabling spatially coherent and high-fidelity animation through linear blend skinning-based Gaussian deformation.
@@
We demonstrate the advantages of our proposed approach through extensive quantitative and qualitative experiments on the challenging ActorsHQ isik2023humanrf and 4D-Dress datasets wang20244ddress . Our method achieves state-of-the-art performance in terms of photometric accuracy, perceptual quality, and articulation realism.
@@ 2 Related Work
One-Shot Human Reconstruction. Reconstructing 3D human avatars from a single image has been explored using both mesh-based and volumetric representations. Mesh-based approaches, including PiFU(-HD) saito2019pifu ; saito2020pifuhd and its extensions ho2024sith ; zhang2023sifu ; alldieck2022phorhum ; sengupta2024diffhuman ; zheng2020pamir ; xiu2023econ ; xiu2023icon ; zhang2023globalcorrelated , typically rely on implicit surfaces or hybrid models shen2021dmtet ; huang2024tech ; yang2024hilo to reconstruct geometry and texture. While some decouple albedo and shading sengupta2024diffhuman ; alldieck2022phorhum ; corona2023s3f , mesh-based methods often struggle to model realistic view-dependent effects. Volumetric representations address these limitations. Works based on NeRFs hu2023sherf ; huang2022elicit ; weng_humannerf_2022_cvpr ; gao2024contex ; buhler2023preface ; buehler2024cafca achieve high fidelity but remain slow to render. Recent methods adopt 3D Gaussian Splatting (3DGS) kerbl20233d for efficient, real-time rendering. Human-focused variants xue2024human3diffusion ; tang2025lgm ; pan2024humansplat ; liu2024humanvdm ; chen2024generalizable use diffusion-generated multiviews voleti2024sv3d to supervise Gaussian reconstruction. However, these typically model static pose-space geometry and lack support for animation under novel poses.
@@
Animatable Avatars. To enable animation, early works like ARCH huang2020arch ; he2021arch++ use canonicalization with LBS deformation, extended by feedforward models hu2023sherf ; peng2024charactergen . These methods, however, depend on accurate body registration, which is often challenging for complex poses and clothing. Concurrent to our work, IDOL zhuang2024idolinstantphotorealistic3d and LHM qiu2025LHM directly predict Gaussian splats on the SMPL-X template to support animation, but model view hallucination and geometry reconstruction with a single model, limiting flexibility. In contrast, decomposes the task into more tractable subproblems of multiview hulluncation, 3D lifting and UV-space mapping. This design enables more robust training, modularity, and better handling of pose- and view-dependent appearance effects essential for animation.
@@
Generative Priors. Generative models such as GANs gan and diffusion models Rombach2022StableDiffusion ; lipmanflow have also been used as strong priors for ill-posed problems such as single-image human reconstruction ho2024sith ; zhang2023sifu ; albahar2023humansgd ; pan2024humansplat ; gao2024contex ; zhang2024humanref ; li2024pshuman . Recent methods apply diffusion models for view synthesis and geometry prediction. Human3Diffusion xue2024human3diffusion jointly trains a 3DGS generator with a multiview diffusion model, but remains limited by low resolution (256×256) and lack of animation support. SiTH ho2024sith and HumanSplat pan2024humansplat synthesize unseen views using single- or multi-view diffusion, while SIFU zhang2023sifu uses a transformer to predict geometry and refines textures with diffusion. Our method builds on both paradigms, i.e., we leverage a pretrained video diffusion model wang2024unianimate to synthesize unseen views of the person, and we use a PatchGAN isola2017image to achieve higher realism and high-frequency details.
@@
Generating 3D Avatars as Latents. Learning generative models of humans in a structured latent space enables controllable synthesis and supports downstream tasks such as few-shot reconstruction, inpainting, and editing. Early 3D methods map latent codes to implicit fields dong2023ag3d ; bergman2022gnarf ; hong2023evad or SDFs noguchi2022unsupervised ; zhang2023getavatar ; deepsdf , often using tri-plane features and adversarial supervision. To improve pose control and spatial detail, recent works incorporate body model priors loper2015smpl ; pavlakos2019smplx and structure latents accordingly hu2024structldm ; chen2023primdiffusion ; abdal2023gaussian ; kolotouros2024avatarpopup ; hong2023evad ; men2024en3d . However, these methods still fall short in texture realism and often rely on optimization-based inversion 10.1145/3658162 ; roich2021pivotal for image/text conditioning. In contrast, our method directly maps an input image to a compact, UV-aligned avatar latent that naturally supports animation, editing, and even avatar interpolation.
@@ 3 Method
Fig. 2 provides a high-level overview of our proposed framework for reconstructing animatable 3D human avatars from a single image. Given the input image $\mathbf{I}_{1}\in\mathbb{R}^{H_{I}\times W_{I}\times C_{I}}$, our method proceeds in three stages: Dream, Lift, and Animate. In the first stage, Dream, we employ a video diffusion model wang2024unianimate to generate plausible multi-view images ${\mathbf{I}_{2}^{n},\dots,\mathbf{I}_{V}^{n}}$ from the input, addressing self-occlusions and incomplete viewpoints. Although visually compelling, these generated views often exhibit inconsistencies. Thus, in the second stage, Lift, we lift these multi-view images into a coherent set of unstructured pose-space 3D Gaussians ${\mathbf{G}_{1}^{p},\dots,\mathbf{G}_{K}^{p}}$ and use a novel transformer-based encoder to transform them into an animation-friendly structured latent representation $\mathbf{Z}$, which is aligned with the UV space of the SMPL-X body model pavlakos2019smplx . Finally, in the third stage, Animate, a Gaussian Parameter Decoder predicts a Gaussian parameter map $\mathbf{F}$ in the UV space from the latent code $\mathbf{Z}$, target pose, and viewpoint conditions. We can then sample structured Gaussians $\mathbf{G}_{1}^{s},\ldots,\mathbf{G}_{N}^{s}$ from the parameter map. These Gaussians can be readily animated via linear blend skinning (LBS) and rendered in real-time: $\mathbf{\hat{I}}=\mathcal{R}(\{\mathbf{G}_{1}^{s},\ldots,\mathbf{G}_{N}^{s}\},\Theta,\mathbf{\pi}),$ where $\mathcal{R}$ is a rendering function that deforms the structured Gaussians $\mathbf{G}_{1}^{s},\ldots,\mathbf{G}_{N}^{s}$ according to target pose $\Theta$ using SMPL-X linear blend skinning, and rasterizes them with camera parameters $\mathbf{\pi}$.
Figure 3: Lifting from multiview images to an avatar latent code. The pose-space reconstruction model produces pixel-aligned Gaussian parameters with corresponding feature maps, denoted pose-space Gaussians and per-Gaussian features. The Gaussians are filtered and subsampled to construct a compact Gaussian feature $\mathbf{X}$ with 2048 Gaussians. These inputs are further processed by a transformer. Specifically, the compact Gaussian feature serves as context (key $\mathbf{K}$ and value $\mathbf{V}$) to a cross-attention layer with queries $\mathbf{Q}$ being the positionally encoded vertex position map $\mathbf{P}$. Finally, the output is reshaped and yields the avatar latent code $\mathbf{Z}$. Please note that this figure omits linear projections, skip-connections, and positional encoding for improved readability.
@@ 3.1 Dream: Multi-view Generation
As illustrated in Fig. 2, our method begins by generating a set of multi-view images consisting of the original input view $\mathbf{I}_{1}$ and a set of novel views $\mathbf{I}_{2}^{n},\ldots,\mathbf{I}_{V}^{n}$. During training, these views are sourced directly from multi-view datasets with known camera calibrations. At inference time, however, we synthesize novel views using a ControlNet-guided variant wang2024unianimate of a video diffusion model wan2025 . Specifically, we first estimate the SMPL-X parameters from the input image he2024magicman and render 2D skeletal poses of the predicted mesh from virtual cameras placed around a 360-degree azimuth. These projected poses serve as control signals to guide the diffusion model in generating photorealistic images from novel viewpoints. While this approach enables hallucination of previously unseen regions and recovers occluded appearance details, the resulting views may still exhibit 3D inconsistencies due to artifacts inherent to the generative diffusion process.
@@ 3.2 Lift: Unstructured Gaussian and Latent Avatar Code Generation
Unstructured Gaussians Reconstruction. We employ a pose-space reconstruction model $\mathcal{G}$ to map the multiview images into a set of pixel-aligned 3D Gaussians. The model uses a U-Net-based architecture augmented with cross-view self-attention, following the design of the Large Gaussian Model (LGM) tang2025lgm . The resulting Gaussians from all views are then fused into a single set of unstructured 3D Gaussians. Reconstructing Gaussians directly in the input pose space is a tractable and effective strategy, as it avoids the highly non-linear mappings required to directly predict avatar geometry in canonical or UV coordinates. This design choice allows our method to faithfully recover rich appearance and geometric details from the input views while remaining robust to inconsistencies.
Latent Avatar Code Generation.
While informative, these unstructured Gaussians are not directly amenable to animation because they lack a consistent topology and structure. To address this, we introduce a transformer-based encoder $\mathcal{F}$ that converts the unstructured Gaussians into a structured latent avatar code $\mathbf{Z}$. The latent code is aligned with the UV space of the SMPL-X model pavlakos2019smplx , which supports expressive deformation and animation via linear blend skinning (LBS). As illustrated in Fig. 3, we first extract per-Gaussian features by concatenating each Gaussian’s raw parameters (e.g., color, position) with intermediate U-Net activations and linearly project them to an embedding space. These features are pixel-aligned by construction, owing to the design of the U-Net. The per-Gaussian features are then filtered and downsampled with farthest-point sampling eldar1997farthest into a more compact Gaussian feature $\mathbf{X}\in\mathbb{R}^{P\times C_{p}}$ where $P$ are the number of filtered Gaussians and $C_{p}$ is the combined dimensionality of the Gaussian parameters and U-Net features.

### v2

@@ Abstract
We introduce Dream, Lift, Animate (DLA), a novel framework that reconstructs animatable 3D human avatars from a single image. This is achieved by leveraging multi-view generation, 3D Gaussian lifting, and pose-aware UV-space mapping of 3D Gaussians. Given an image, we first dream plausible multi-views using a video diffusion model, capturing rich geometric and appearance details. These views are then lifted into unstructured 3D Gaussians. To enable animation, we propose a transformer-based encoder that models global spatial relationships and projects these Gaussians into a structured latent representation aligned with the UV space of a parametric body model. This latent code is decoded into UV-space Gaussians that can be animated via body-driven deformation and rendered conditioned on pose and viewpoint. By anchoring Gaussians to the UV manifold, our method ensures consistency during animation while preserving fine visual details. DLA enables real-time rendering and intuitive editing without requiring post-processing. Our method outperforms state-of-the-art approaches on the ActorsHQ and 4D-Dress datasets in both perceptual quality and photometric accuracy. By combining the generative strengths of video diffusion models with a pose-aware UV-space Gaussian mapping, DLA bridges the gap between unstructured 3D representations and high-fidelity, animation-ready avatars.
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
@@
Figure 1: We propose Dream, Lift, Animate, a novel framework to reconstruct high-fidelity, animatable 3D human avatars from a single image by generating multi-view images, lifting them to 3D Gaussians, and mapping them to a pose-aware UV space (Fig. 2). Our approach enables realistic animation and outperforms prior methods in visual quality (Fig. 5 and Tbl. 1). Watch videos and more at https://research.nvidia.com/labs/dair/dream-lift-animate.
@@ 1 Introduction
Figure 2: Overview of the proposed Dream, Lift, Animate (DLA) framework for reconstructing animatable 3D human avatars from a single image. In the Dream stage (Sec. 3.1), we synthesize novel views from the input using a diffusion-based generator. In the Lift stage (Sec. 3.2 and Fig. 3), we project the multi-view images into a set of unstructured 3D Gaussians in the pose space using a learned Gaussian reconstruction model $\mathcal{G}$. Subsequently, we learn a transformer encoder $\mathcal{F}$ to map 3D Gaussians to a structured latent code $\mathbf{Z}$ in the UV space of a parametric body model. In the Animate stage (Sec. 4 and Fig. 4), we decode the avatar code into a pose- and view-aware Gaussian parameter map $\mathbf{F}$. This structured representation enables realistic animation and rendering via deformation with a body model.
@@
Existing methods typically fall short in simultaneously addressing these challenges. Recent approaches [20, 69, 18] leverage the powerful generative capabilities of multi-view diffusion models [51] to hallucinate missing appearance information; however, these models inherently produce inconsistencies across generated views, leading to avatars that appear overly smoothed, lack critical details, or introduce visually jarring artifacts when animated. Conversely, methods relying on template-based rigging approaches, such as automatically transferring skinning weights from canonical human body models (e.g., SMPL [39]), typically require careful fitting procedures [77, 24, 25, 57]. While powerful, these methods can still encounter challenges when handling non-standard poses, complex clothing, and significant self-occlusions frequently encountered in unconstrained, real-world scenarios, thus affecting their robustness and broader applicability.
@@
To overcome these limitations, we propose Dream, Lift, Animate (DLA), a novel framework designed specifically for reconstructing high-quality, animatable 3D human avatars from a single image. Our method addresses the core reconstruction problem by decomposing it into three complementary stages, each carefully designed to handle limitations inherent in the preceding steps. Specifically, we first leverage a pretrained video diffusion model [64] to Dream plausible multi-view observations of the subject, effectively hallucinating realistic geometric and appearance details even in regions unseen from the input viewpoint. Despite the significant progress made by recent diffusion-based approaches, their inherent view-to-view inconsistencies prevent direct avatar reconstruction [69]. To resolve this, we next Lift these generated views into an intermediate, unstructured 3D representation based on 3D Gaussian primitives [32] in the input pose space, effectively aggregating appearance and geometry cues from the multi-view observations. Subsequently, to enable animation, we map these unstructured Gaussians into a structured UV-space representation. For this, we use a transformer-based encoder [72] that models global spatial relationships among the unstructured 3D Gaussians and projects them into a structured latent avatar representation, which aligns with the UV space of the SMPL-X body model [46]. This learned mapping also allows our model to reconcile inconsistencies introduced during the initial multi-view generation seamlessly. Finally, in the Animate stage, we present a pose- and view-aware Gaussian parameter decoder that converts this latent avatar representation into a UV-space map of Gaussian parameters, enabling spatially coherent and high-fidelity animation through linear blend skinning-based Gaussian deformation.
@@
We demonstrate the advantages of our proposed approach through extensive experiments on the challenging ActorsHQ [28], 4D-Dress  [63], and SHHQ [13] datasets. Our method achieves state-of-the-art performance in terms of photometric accuracy, perceptual quality, and articulation realism.
@@ 2 Related Work
One-Shot Human Reconstruction. Reconstructing 3D human avatars from a single image has been explored using both mesh-based and volumetric representations. Mesh-based approaches, including PiFU(-HD) [52, 53] and its extensions [20, 77, 3, 54, 78, 68, 67, 76], typically rely on implicit surfaces or hybrid models [55, 25, 70] to reconstruct geometry and texture. While some decouple albedo and shading [54, 3, 10], mesh-based methods often struggle to model realistic view-dependent effects. Volumetric representations address these limitations. Works based on NeRFs [22, 24, 65, 14, 6, 7] achieve high fidelity but remain slow to render. Recent methods adopt 3D Gaussian Splatting (3DGS) [32] for efficient, real-time rendering. Human-focused variants [69, 58, 43, 38, 8] use diffusion-generated multiviews [61] to supervise Gaussian reconstruction. However, these typically model static pose-space geometry and lack support for animation under novel poses.
@@
Animatable Avatars. To enable animation, early works like ARCH [27, 17] use canonicalization with LBS deformation, extended by feedforward models [22, 47]. These methods, however, depend on accurate body registration, which is often challenging for complex poses and clothing. IDOL [80] and LHM [49] (concurrent work) directly predict Gaussian splats on the SMPL-X template to support animation. However, their deformation module does not consider pose-dependent effects. In contrast, our approach conditions the Gaussian parameters on the target pose, enabling pose-dependent effects. In addition, our approach is substantially more lightweight because it decomposes the task into more tractable subproblems of multiview hallucination, 3D lifting, and UV-space mapping. While IDOL and LHM train on multiple nodes (32 NVIDIA A100/H100 GPUs), our model only requires a single node (8 NVIDIA A100 GPUs). Our lightweight design enables modularity and better handling of pose- and view-dependent appearance effects essential for animation.
@@
Generative Priors. Generative models such as GANs [16] and diffusion models [51, 37] have also been used as strong priors for ill-posed problems such as single-image human reconstruction [20, 77, 2, 43, 14, 73, 36, 40, 26]. Recent methods apply diffusion models for view synthesis and geometry prediction. Human3Diffusion [69] jointly trains a 3DGS generator with a multiview diffusion model, but remains limited by low resolution ($256\times 256$) and a lack of animation support. SiTH [20] and HumanSplat [43] synthesize unseen views using single- or multi-view diffusion, while SIFU [77] uses a transformer to predict geometry and refines textures with diffusion. Our method builds on both paradigms, i.e., we leverage a pretrained video diffusion model [64] to synthesize unseen views of the person, and we use a PatchGAN [29] to achieve higher realism and high-frequency details.
@@
Generating 3D Avatars as Latents. Learning generative models of humans in a structured latent space enables controllable synthesis and supports downstream tasks such as few-shot reconstruction, inpainting, and editing. Early 3D methods map latent codes to implicit fields [11, 4, 21] or SDFs [42, 75, 44], often using tri-plane features and adversarial supervision. To improve pose control and spatial detail, recent works incorporate body priors [39, 46] with structured latents [23, 9, 1, 33, 21, 41]. However, these methods still fall short in texture realism and often rely on optimization-based inversion [66, 50] for image/text conditioning. In contrast, our method directly maps an input image to a compact, UV-aligned avatar latent that naturally supports animation, editing, and even interpolation.
@@ 3 Method
Fig. 2 provides a high-level overview of our proposed framework for reconstructing animatable 3D human avatars from a single image. Given the input image $\mathbf{I}_{1}\in\mathbb{R}^{H_{I}\times W_{I}\times C_{I}}$, our method proceeds in three stages: Dream, Lift, Animate. In the first stage, Dream, we employ a video diffusion model [64] to generate plausible multi-view images $\mathbf{I}_{2}^{n},\dots,\mathbf{I}_{V}^{n}$} from the input, addressing self-occlusions and incomplete viewpoints. Although visually compelling, these generated views often exhibit inconsistencies. Thus, in the second stage, Lift, we lift these multi-view images into a coherent set of unstructured pose-space 3D Gaussians $\mathbf{G}_{1}^{p},\dots,\mathbf{G}_{K}^{p}$} and use a novel transformer-based encoder to transform them into an animation-friendly structured latent representation $\mathbf{Z}$, which is aligned with the UV space of the SMPL-X body model [46]. Finally, in the third stage, Animate, a Gaussian Parameter Decoder predicts a Gaussian parameter map $\mathbf{F}$ in the UV space from the latent code $\mathbf{Z}$, target pose, and viewpoint conditions. We can then sample structured Gaussians $\mathbf{G}_{1}^{s},\ldots,\mathbf{G}_{N}^{s}$ from the parameter map. These Gaussians can be readily animated via linear blend skinning (LBS) and rendered in real-time: $\mathbf{\hat{I}}=\mathcal{R}(\{\mathbf{G}_{1}^{s},\ldots,\mathbf{G}_{N}^{s}\},\Theta,\mathbf{\pi}),$ where $\mathcal{R}$ is a rendering function that deforms the structured Gaussians $\mathbf{G}_{1}^{s},\ldots,\mathbf{G}_{N}^{s}$ according to target pose $\Theta$ using SMPL-X linear blend skinning, and rasterizes them with camera parameters $\mathbf{\pi}$.
Figure 3: Lift from multiview images to an avatar latent code. The pose-space reconstruction model produces pixel-aligned Gaussian parameters with corresponding feature maps, denoted pose-space Gaussians and per-Gaussian features. The Gaussians are filtered and subsampled to construct a compact Gaussian feature $\mathbf{X}$ with 2048 Gaussians. These inputs are further processed by a transformer. Specifically, the compact Gaussian feature serves as context (key $\mathbf{K}$ and value $\mathbf{V}$) to a cross-attention layer with queries $\mathbf{Q}$ being the positionally encoded vertex position map $\mathbf{P}$. Finally, the output is reshaped and yields the avatar latent code $\mathbf{Z}$. This figure omits linear projections, skip-connections, and positional encoding for improved readability.
@@ 3.1 Dream: Multi-view Generation
As illustrated in Fig. 2, our method begins by generating a set of multi-view images consisting of the original input view $\mathbf{I}_{1}$ and a set of novel views $\mathbf{I}_{2}^{n},\ldots,\mathbf{I}_{V}^{n}$. During training, these views are sourced directly from multi-view datasets with known camera calibrations. At inference time, however, we synthesize novel views using a ControlNet-guided variant [64] of a video diffusion model [62]. Specifically, we first estimate the SMPL-X parameters from the input image [18] and render 2D skeletal poses of the predicted mesh from virtual cameras placed around a 360-degree azimuth. These projected poses serve as control signals to guide the diffusion model in generating photorealistic images from novel viewpoints. While this approach enables hallucination of previously unseen regions and recovers occluded appearance details, the resulting views may still exhibit 3D inconsistencies due to artifacts inherent to the generative diffusion process. Please see the supplementary material for more details.
@@ 3.2 Lift: Unstructured Gaussian and Latent Avatar Code Generation
Unstructured Gaussians Reconstruction. We employ a pose-space reconstruction model $\mathcal{G}$ to map the multiview images into a set of pixel-aligned 3D Gaussians. The model uses a U-Net-based architecture augmented with cross-view self-attention, following the design of the Large Gaussian Model (LGM) [58]. The resulting Gaussians from all views are then fused into a single set of unstructured 3D Gaussians. Reconstructing Gaussians directly in the input pose space is a tractable and effective strategy, as it avoids the highly non-linear mappings required to directly predict avatar geometry in canonical or UV coordinates. This design choice allows our method to faithfully recover rich appearance and geometric details from the input views while remaining robust to inconsistencies.
Latent Avatar Code Generation. While informative, these unstructured Gaussians are not directly amenable to animation because they lack a consistent topology and structure. To address this, we introduce a transformer-based encoder $\mathcal{F}$ that converts the unstructured Gaussians into a structured latent avatar code $\mathbf{Z}$. The latent code is aligned with the UV space of the SMPL-X model [46], which supports expressive deformation and animation via linear blend skinning (LBS). As illustrated in Fig. 3, we first extract per-Gaussian features by concatenating each Gaussian’s raw parameters (e.g., color, position) with intermediate U-Net activations and linearly project them to an embedding space. These features are pixel-aligned by construction, owing to the design of the U-Net. The per-Gaussian features are then filtered and downsampled with farthest-point sampling [12] into a more compact Gaussian feature $\mathbf{X}\in\mathbb{R}^{P\times C_{p}}$ where $P$ are the number of filtered Gaussians and $C_{p}$ is the combined dimensionality of the Gaussian parameters and U-Net features.